OpenStudy (anonymous):

Can some body give me reassurance that I'm doing this problem correctly: I'm finding the to a right triangle that goes like this= a is 2.48m , b is 1.45m what would be the value for c? Do I multiply a & b or do I squar them then multiply, help?

OpenStudy (anonymous):

square them and then multiply

OpenStudy (anonymous):

like if youre doinh the pythagorean theorem

OpenStudy (tkhunny):

c is the side opposite the right angle and is always the side of greatest length. The application of the Pythagorean Theorem solves the problem. \(c = \sqrt{a^{2}+b^{2}}\) Let's see what you get.

OpenStudy (anonymous):

is the first thing I should do is find the powers for a & b first, next should I square them off, finally subtract a from b to give me c, then I square c for my final answer?

OpenStudy (tkhunny):

Why are you subtracting anything? When finding the side of greatest length, there will be no subtraction.

OpenStudy (anonymous):

I'm so very sorry I was thinking about my bank account, all ways in the negative. lol I meant add.

OpenStudy (tkhunny):

Why wld you square 'c'? You don't know what it is. How can you square it. Square a Square b Add those two results. Find the square root of that result. Done.

OpenStudy (anonymous):

so I have 8.36, now do I square that?

OpenStudy (anonymous):

2.90??????????

OpenStudy (anonymous):

I'm sorry for the confusion, I know b & c i'm trying to find a.

OpenStudy (anonymous):

I just had an epiphany, so do I take b^ from c^ for a^?

OpenStudy (anonymous):

once I have a^ is that the correct answer or should I square it?

OpenStudy (tkhunny):

You may be in the wrong class. You are struggling far too much with the notation. Square a 2.48m ==> (2.48m)^2 = 6.1504m^2 Square b 1.45m ==> (1.45)^2 = 2.1025m^2 Add those two results. ==> 6.1504m^2 + 2.1025m^2 = 8.2529m^2 Find the Square Root of that Result \(\sqrt{8.2529m^2} = 2.87m\)